# Attacking LLM Watermarks by Exploiting Their Strengths

### v1

@@ 4.1 Evaluation
Observation #1 Robust watermarks are inherently susceptible to spoofing attacks.
Experiment Setup. We assess the effectiveness of the spoofing attack exploiting the robustness property by determining the maximum portion of toxic tokens that can be inserted into the watermarked text without altering the watermark detection result. Specifically, we generate a list of 200200200 toxic tokens and insert them at random positions of the watermarked outputs. We utilize 500500500 prompts data from OpenGen [KSK+23] dataset and query the watermarked language models (including Llama-2-7B [TMS+23] and OPT-1.3B [ZRG+22]) to generate the watermarked outputs. We evaluate three SOTA watermarks including KGW [KGW+23a], Unigram [ZALW24], and Exp [KTHL23]. We choose the default watermarking hyperparameters. In our experiments, we default to a maximum of 200 new tokens for KGW and Unigram, and 70 for Exp, due to its complexity in the watermark detection. 70 is also the maximum number of tokens the authors of Exp evaluated in their paper [KTHL23].
Evaluation Result. We report the maximum portion of the inserted toxic tokens relative to the original watermarked sentence length in Fig. 1. We also present the confidence of the OpenAI moderation model [Ope23b] in identifying the content as violating their usage policy [Ope23c] due to the inserted toxic tokens in Fig. 1.
[…]
Figure 1: Spoofing results. The LHS shows the maximum portion of toxic tokens permitted for insertion. The RHS shows the OpenAI moderation’s confidence in detecting the toxic content.
@@ 4.2 Discussion
Guideline #1 Robust watermarks may need to compromise on robustness to mitigate the possibility of spoofing attacks.
@@
Spoofing attacks are easy to execute in practice. We are not aware of any existing watermarks that consider such attacks during the design and deployment, and existing robust watermarks are inherently vulnerable to such attacks. In particular, we highlight the tension that exists between watermark robustness and spoofing feasibility: While enhancing robustness makes it more difficult to remove watermarks from edited sentences, this feature can be directly exploited as an attack vector, where the attacker might edit the watermarked content to include harmful text while ensuring that the output as a whole remains watermarked. We deem that this attack is challenging to defend against, especially considering the spoofing example in Table 1, where by only editing a single token, the entire content becomes incorrect. As we show through our feasibility analysis (Theorem 1) it may be hard, if not impossible, in practical settings to detect whether a particular token is from the attacker by using robust watermark detection algorithms. Thus, practitioners should carefully weigh the risks of removal vs. spoofing attacks for the model at hand, and consider reducing watermark robustness to mitigate the possibility of spoofing. The general public should also be aware of this pitfall—i.e., that it is currently quite easy to generate content which appears to be watermarked but is in fact not generated by the underlying model in question.
5 Attacking Quality-Preserving Watermarks
@@
SOTA watermarking schemes [KGW+23a, FGJ+23, CGZ23, KTHL23, ZALW24, KGW+23b] aim to ensure the watermarked text retains its high quality by maintaining an “unbiasedness” property. Formally:
@@
$\mathbb{E}_{sk\in\mathcal{S}}(M_{\text{wm}}(\textbf{x},sk))\approx_{\epsilon}M_{\text{orig}}(\textbf{x}),$ (5)
@@
the expected distribution of watermarked output over the watermark key space $sk\in\mathcal{S}$ is close to the output distribution without a watermark, differing by a distance of $\epsilon$. We note that Exp [KTHL23] is rigorously unbiased, and KGW [KGW+23a] and Unigram [ZALW24] slightly shift the watermarked distributions during watermark embedding.
As a common practice in cryptography and also suggested by KGW [KGW+23a] or inherent in prior watermarks (e.g., Exp [KTHL23]), the LLM service provider may use multiple watermark keys to enhance security against brute-force attacks in distinguishing the watermarked tokens [SKB+23]. However, we demonstrate that using multiple keys can potentially introduce new vulnerabilities and allow malicious users to remove watermarks with only a few queries to the watermarked LLM.
@@
Here, the intuition is that, given the quality-preserving nature, malicious users can estimate the distribution of the LLM without any watermark by making multiple queries to the watermarked LLM under different keys for each token. As this attack estimates the original, unwatermarked distribution, the quality of the generated content is preserved.
@@
Attack Procedure. An attacker queries a watermarked model with an input x of length $t$ under $n$ different watermark keys, observing $n$ subsequent tokens $\textbf{x}_{t+1}$. They then create a frequency histogram of these tokens and sample the most frequent one. This sampled token matches the result of greedy sampling on an unwatermarked output distribution with a nontrivial probability. Consequently, the attacker can progressively eliminate watermarks while maintaining a high quality of the synthesized content.
Key/Query Number Analysis. Now we analyze the number of required queries under different keys to estimate the token with the highest probability without a watermark. We have the following probability bound for KGW and Unigram (proof deferred to Appendix D), and present the bound for Exp in Appendix E.
Theorem 2 (Probability bound of unwatermarked token estimation).
$1-\sum_{k=0}^{\lfloor n/2\rfloor}\binom{n}{k}\gamma^{k}(1-\gamma)^{n-k}\times p(k),$ (6)
[…]
@@ 5.1 Evaluation
Observation #2 While increasing the number of watermark keys enhances the expected output quality and security against brute-force attacks, it also increases vulnerability to watermark-removal attacks.
(a) Z-Score.
(b) Perplexity.
Figure 2: Watermark-removal on KGW watermark and Llama-2-7B model with different number of watermark keys $n$. Higher z-score reflects more confidence in watermarking and lower perplexity indicates better sentence quality.
Experiment Setup. Similar to the evaluation setup in Sec. 4.1, we evaluate three SOTA watermarks, KGW, Unigram, and Exp on two models, Llama-2-7B and Opt-1.3B, and OpenGen dataset. We evaluate the detection scores (z-score or p-value) and the output perplexity (PPL) evaluated using GPT3 [OWJ+22] (see the introduction of perplexity in Appendix A). We use the default watermark hyperparameters.
Evaluation Result. As shown in Fig. 2(a), we can significantly reduce the detection confidence scores by using more keys. We also present the median detection scores of the unwatermarked content using different numbers of keys in Fig. 2(a). Since the detection algorithm returns the highest z-score among all the keys [KGW+23a, KTHL23], the expected z-scores of unwatermarked content become higher if using more keys. The scores of the watermark-removal closely resemble those of unwatermarked content detection results when we use more than 7 keys. In essence, to remove the watermark, an attacker only needs to query the watermarked LLM 7 times for each token under different keys. Fig. 2(b) suggests that using more keys improves the output content quality. This is because, with a greater number of keys, there’s a higher probability for an attacker to accurately estimate the unwatermarked distribution, which is consistent with our analysis in Eq. 6. We observe that in practice, 7 keys suffice to produce high-quality content comparable to the unwatermarked content. This observation remains consistent across various watermarking schemes and models, we defer more results to Appendix G.
@@ 5.2 Discussion
Guideline #2 Using fewer keys, while potentially leading to reduced quality and security against brute-force attacks, may be necessary to mitigate watermark-removal attacks.
Many existing watermarking schemes recommend using multiple keys to ensure distortion-free watermarks and enhance security against brute-force attacks. However, we reveal that a conflict exists between maintaining quality/security and the feasibility of removing watermarks. We suggest using fewer keys and compromising slightly on quality and security against brute-force attacks in practice to achieve a balanced trade-off. In our experiments, we observed that limiting the use of keys to no more than three can impede the success of watermark-removal attacks and offer reasonable resistance to brute-force attacks. However, it’s important to acknowledge the need to compromise the quality of the Exp watermark [KTHL23] given the potency of the watermark-removal attack. Additionally, as a simple defense, we recommend that LLM service providers limit their query rates for users, particularly for users who have been identified as being potentially malicious.
@@ 6 Attacking Watermark Detection APIs
Another common feature of modern LLM watermarks is corresponding public detection APIs, which allow the general public to easily verify if a passage of text is AI-generated [FGJ+23, KGW+23a, SBC+19, MLK+23]. However, this property can also be exploited to launch watermark-removal and spoofing attacks. In the following sections, we first introduce the attack procedures and then propose suggestions and defenses to mitigate these attacks.
@@ 6.1 Attack Procedures
Watermark-Removal Attack. For the watermark-removal attack, we consider an attacker who has access to the target watermarked LLM API, and can query the watermark detection results. The attacker feeds a prompt into the watermarked LLM, which generates the response in an auto-regressive manner. For the token $\textbf{x}_{i}$ the attacker will generate a list of possible replacements for $\textbf{x}_{i}$. This list can be generated by querying the watermarked LLM, querying the local model, or simply returned by the watermarked LLM. In this work, we choose the third approach because of its simplicity and guarantee of synthesized sentences’ quality. This is a common assumption made by prior works [NKIH23], and such an API is also provided by OpenAI ($\mathrm{top\_logprobs=5}$). Additionally, returning the top logprob values can benefit the normal users in understanding the model confidence, debugging and analyzing the model’s behavior, customizing sampling strategies, etc.
Consider that the top $L=5$ tokens and their probabilities are returned to the attackers. The probability that the attacker can find an unwatermarked token in the token candidates’ list of length $L$ is $1-\gamma^{L}$ for KGW and Unigram, which becomes sufficiently large given $L=5$ and $\gamma=0.5$.
The attacker will query the detection using these replacements and sample a token based on their probabilities and detection scores to remove the watermark while preserving a high output quality. See Alg. 1 in Appendix F.
Spoofing Attack. Spoofing attacks follow a similar procedure where the attacker can generate (harmful) content using a local model, and select tokens that yield higher confidence scores upon watermark detection queries. Thanks to the robustness of the LLM watermarks, the attackers don’t need to ensure every single token carries a watermark; only that the overall detection confidence score surpasses the threshold, thereby treating synthesized content as if generated by the watermarked LLM. Please refer to Alg. 2 in Appendix F for the detailed algorithm.
(a) Z-Score/P-Value of removal.
(b) Perplexity of removal.
@@ 6.2 Evaluation
Observation #3 Public detection APIs can be exploited to launch both watermark-removal and spoofing attacks.
Experiment Setup. We use the same evaluation setup as Sec. 4.1 and Sec. 5.1. We evaluate the feasibility of the attacks exploiting the detection API on three watermarks (KGW, Unigram, and Exp), two LLMs (Llama-2-13B, OPT-1.3B), and OpenGen dataset. We evaluate the detection scores for both the watermark-removal and the spoofing attacks. We also report the number of queries to the detection API. Furthermore, for the watermark-removal attack, where the attackers care more about the output quality, we report the output PPL. For spoofing attacks, the attackers’ local models are Llama-2-7B and OPT-1.3B.
Evaluation Result. The results are shown in Fig. 3. Watermark-removal attack exploiting the detection API significantly reduces detection confidence while maintaining high output quality as shown in Fig. 3(a) and Fig. 3(b). For instance, for the KGW watermark on Llama-2-7B model, we achieve a median z-score of 1.431.431.43, which is much lower than the threshold 444. The PPL is also close to the watermarked outputs (6.176.176.17 vs. 6.286.286.28). We observe that the Exp watermark has a higher PPL than the other two watermarks. This is because that Exp watermark is deterministic, while other watermarks enable random sampling during inference. Our attack also employs sampling based on the token probabilities and detection scores, thus we can improve the output quality for the Exp watermark.
The spoofing attacks also significantly boost the detection confidence even though the content is not from the watermarked LLM, as depicted in Fig. 3(c). We report the attack success rate (ASR) and the number of queries for both of the attacks in Table 2. The ASR quantifies how much of the generated content surpasses or falls short of the detection threshold. These attacks use a reasonable number of queries to the detection API and achieve a high success rate, demonstrating practical feasibility. We observe consistent results on OPT-1.3B model, please see Appendix H.

### v3

@@ 4.1 Evaluation
Experiment Setup.  We assess the effectiveness of our piggyback spoofing attack by using the two editing strategies discussed above. Through toxic token insertion, we study the limits of how many tokens can be inserted into the watermarked content. Using fluent inaccurate editing, we show that piggyback spoofing can generate fluent, watermarked, but inaccurate results at scale. Specifically, for the toxic token insertion, we generate a list of $200$ toxic tokens and insert them at random positions in the watermarked output. For the fluent inaccurate editing, we edit the watermarked sentence by querying GPT4 using the prompt “Modify less than 3 words in the following sentence and make it inaccurate or have opposite meanings.” Unless otherwise specified, in the evaluations of this work, we utilize $500$ prompts data from OpenGen [19] dataset, and query the watermarked language models (LLAMA-2-7B [38] and OPT-1.3B [45]) to generate the watermarked outputs. We evaluate three SOTA watermarks including KGW [17], Unigram [46], and Exp [20], using the default watermarking hyperparameters. In our experiments, we default to a maximum of 200 new tokens for KGW and Unigram, and 70 for Exp, due to its complexity in the watermark detection. 70 is also the maximum number of tokens the authors of Exp evaluated in their paper [20].
(a) Toxic token insertion.
(b) Fluent inaccurate editing.
Figure 1: Piggyback spoofing of robust watermarks. (a) We can insert a large number of toxic tokens in robustly watermarked text without changing the watermark detection result, resulting in text that is likely to be identified as toxic. (b) We can use GPT4 to automatically modify watermarked text, making it appear inaccurate while retaining fluency.
Evaluation Result. We report the maximum portion of the inserted toxic tokens relative to the original watermarked sentence length on LLAMA-2-7B model in Fig. 1. We also present the confidence of the OpenAI moderation model [30] in identifying the content as violating their usage policy [31] due to the inserted toxic tokens in Fig. 1. Our findings show that we can insert a significant number of toxic tokens into content generated by all the robust watermarking schemes, with a median portion higher than $20\%$, i.e., for a $200$-token sentence, the attacker can insert a median of $40$ toxic tokens into it. These toxic sentences are then identified as violating OpenAI policy rules with high confidence scores, whose median is higher than 0.8 for all the watermarking schemes we study. The average confidence scores for content before attack are around 0.01. The empirical data on the maximum portion of inserted toxic tokens aligns with our analysis in Appendix D. We further validate this analysis in Fig. 3(e) of Appendix E, showing that attackers can insert nontrivial portions of toxic tokens into the watermarked text to launch piggyback spoofing attacks. Notably, the more robust the watermark is, the more tokens can effectively be inserted. We present the results on OPT-1.3B in Appendix G.
In Fig. 1, we report the PPL and watermark detection scores of the piggyback results on KGW and LLAMA-2-7B by the fluent inaccurate editing strategy. We show that we can successfully generate fluent results, with a slightly higher PPL. $94.17\%$ of the piggyback results have a z-score higher than the default threshold $4$. We randomly sample $100$ piggyback results and manually check that most of them ($92\%$) are fluent and have inaccurate or opposite content from the original watermarked content. See examples in Appendix F. The results show that we can generate watermarked, fluent, but inaccurate content at scale with an ASR higher than 90%.
@@ 4.2 Discussion
Guideline #1 Robust watermarks are inherently vulnerable to spoofing attacks and are not suitable as proof of content authenticity alone. To mitigate spoofing while preserving robustness, it may be necessary to combine additional measures such as signature-based fragile watermarks.
@@
Our results highlight that piggyback spoofing attacks are easy to execute in practice. LLM watermarks typically do not consider such attacks during design and deployment, and existing robust watermarks are inherently vulnerable to such attacks. We highlight the contradiction between the watermark robustness and the piggyback spoofing feasibility. We consider this attack to be challenging to defend against, especially considering examples such as those in Table 1 and Appendix F, where by only editing a single token, the entire content becomes incorrect. It is hard, if not impossible, to detect whether a particular token is from the attacker by using robust watermark detection algorithms. Thus, practitioners should weigh the risks of removal vs. piggyback spoofing attacks for the model at hand. A feasible strategy to mitigate spoofing attacks is by requiring proof of digital signatures on the LLM generated content. However, while an attacker without access to the private key cannot spoof, it is worth nothing that this strategy is still vulnerable to watermark-removal attacks, as a single editing can invalidate the original signature.
5 Attacking Stealing-Resistant Watermarks
As discussed in Sec. 2, many works have explored the possibility of launching watermark stealing attacks to infer the secret pattern of the watermark, which can then enable spoofing and removal attacks [35, 15, 11]. A natural and effective defense against watermark stealing is using multiple watermark keys during embedding, which is a common practice in cryptography and also suggested by prior watermarks and work in watermark stealing [17, 20, 15]. Unfortunately, we demonstrate that using multiple keys can in turn introduce new watermark-removal attacks.
@@ 5 Attacking Stealing-Resistant Watermarks
In particular, SOTA watermarking schemes [17, 10, 6, 20, 46, 18] aim to ensure the watermarked text retains its high quality and the private watermark patterns are not easily distinguished by maintaining an “unbiasedness” property:
@@
$\footnotesize\mathbb{E}_{sk\in\mathcal{S}}(M_{\text{wm}}(\textbf{x},sk))\approx_{\epsilon}M_{\text{orig}}(\textbf{x}),$ (4)
@@
i.e., the expected distribution of watermarked output over the watermark key space $sk\in\mathcal{S}$ is close to the output distribution without a watermark, differing by a distance of $\epsilon$. We note that Exp [20] is “distortion free” for a single text sample, and KGW [17] and Unigram [46] slightly shift the watermarked distributions. We note that stealing attacks won’t work on rigorously unbiased watermarks.
@@
The insight of our proposed watermark-removal attack is that given the “unbiasedness” nature of watermarks and considering multiple keys may be used during watermark embedding, malicious users can estimate the output distribution without any watermark by querying the watermarked LLM multiple times using the same prompt. As this attack estimates the original, unwatermarked distribution, the quality of the generated content is preserved.
@@
Attack Procedure. An attacker queries a watermarked model with an input x multiple times, observing $n$ subsequent tokens $\textbf{x}_{t+1}$. This is easy for text completion model APIs, and chat model APIs can also be easily attacked by constructing a prompt to ask the chat model to complete a partial sentence without any prefix. The attacker then creates a frequency histogram of these tokens and samples according to the frequency. This sampled token matches the result of sampling on an unwatermarked output distribution with a nontrivial probability. Consequently, the attacker can progressively eliminate watermarks while maintaining a high quality of the synthesized content. We present a formal analysis of the number of required queries in Appendix H.
@@ 5.1 Evaluation
Experiment Setup.  Our watermarks, models and datasets settings are the same as Sec. 4.1. We study the trade-off between resistance against watermark stealing and watermark-removal attacks by evaluating a recent watermark stealing attack [15]. In this attack, we query the watermarked LLM to obtain 2.2 million tokens in total to estimate the watermark pattern and then launch spoofing attacks using the estimated watermark pattern. We follow their assumptions that the attacker can access the unwatermarked tokens’ distribution. In our watermark removal attack, we consider that the attacker has observations with different keys. We evaluate the detection scores (z-score or p-value) and the output perplexity (PPL, evaluated using GPT3 [32]). The detection algorithm returns the maximum detection score across all the keys, which increases the expectation of unwatermarked detection results. Thus, we set the detection thresholds for different keys to keep the false positive rates (FPR) below 1e-3 and report the attack success rates (ASR). We use default watermark hyperparameters.
Evaluation Result. As shown in Fig. 2(a), using multiple keys can effectively defend against watermark stealing attacks. With a single key, the ASR is $91\%$, which matches the results reported in [15]. We observe that using three keys can effectively reduce the ASR to $13\%$, and using more than 7 keys, the ASR of the watermark stealing is close to zero. However, using more keys also makes the system vulnerable to our watermark-removal attacks as shown in Fig. 2(b). When we use more than $7$ keys, the detection scores of the content produced by our watermark removal attacks closely resemble those of unwatermarked content and are much lower than the detection thresholds, with ASRs higher than $97\%$. Fig. 2(c) suggests that using more keys improves the quality of the output content. This is because, with a greater number of keys, there is a higher probability for an attacker to accurately estimate the unwatermarked distribution, which is consistent with our analysis in Appendix H. We observe that in practice, 7 keys suffice to produce high-quality content comparable to the unwatermarked content. These observations remain consistent across various watermarking schemes and models; for additional results see Appendix J. We note that the numbers are not exactly the same as [15], as we consider a more realistic attacker with less queries to the watermarked LLM.
(a) Z-Score and attack success rate (ASR) of watermark stealing [15].
(b) Z-Score and attack success rate (ASR) of watermark-removal.
(c) Perplexity (PPL) of watermark-removal.
Figure 2: Spoofing attack based on watermark stealing [15] and watermark-removal attacks on KGW watermark and LLAMA-2-7B model with different number of watermark keys $n$. Higher z-score reflects more confidence in watermarking and lower perplexity indicates better sentence quality. The attack success rates are based on the threshold with FPR@1e-3.
@@ 5.2 Discussion
Guideline #2 Using a larger number of watermarking keys can defend against watermark stealing attacks, but increases vulnerability to watermark-removal attacks. Limiting users’ query rates can help to mitigate both attacks.
Many prior works have suggested using multiple keys to defend against watermark stealing attacks. However, in this study, we reveal that a conflict exists between improving resistance to watermark stealing and the feasibility of removing watermarks. Our evaluation results show that finding a "sweet spot" in terms of the number of keys to use to mitigate both the watermark stealing and the watermark-removal attacks is not trivial. For example, our watermark-removal attack achieves a high ASR of $36.2\%$ just using three keys, and the corresponding watermark stealing-based spoofing’s ASR is $13.0\%$. Using more keys can decrease the watermark stealing-based spoofing’s ASR, but at the cost of making the system more vulnerable to watermark removal and vice-versa. We note that the ASRs with three keys are not negligible, thus limiting the ability of potentially malicious users is necessary in practice to mitigate these attacks. As a practical defense, we evaluate watermark stealing with various query limits on the watermarked LLM, and found that the ASR can be significantly reduced by limiting the attacker’s query rate. Detailed results can be found in Appendix J. Given the trade-off that exists, we suggest that LLM service providers consider “defense-in-depth” techniques such as anomaly detection, query rate limiting, and user identification verification.
@@ 6 Attacking Watermark Detection APIs
It is still an open question whether watermark detection APIs should be made publicly available to users. Although this makes it easier to detect watermarked text, it is a commonly acknowledged that it will make the system vulnerable to attacks [1] given the existence of oracle attacks [3, 7, 23, 16]. Here, we study this statement more precisely by examining the specific risk trade-offs that exist, as well as introducing a novel defense that may make the public detection API more feasible in practice. In the following sections, we first introduce attacks that exploit the APIs and then propose suggestions and defenses to mitigate these attacks.
@@ 6.1 Attack Procedures
Watermark-Removal Attack. For the watermark-removal attack, we consider an attacker who has access to the target watermarked LLM’s API, and can query the watermark detection results. The attacker feeds a prompt into the watermarked LLM, which generates the response in an auto-regressive manner, similar to how LLMs generate sentences. That is, the attacker will select each token based on the prior tokens and the detection results. For the token $\textbf{x}_{i}$ the attacker will generate a list of possible replacements for $\textbf{x}_{i}$. This list can be generated by querying the watermarked LLM, querying a local model, or simply returned by the watermarked LLM. In this work, we choose the third approach because of its simplicity and guarantee of synthesized sentences’ quality. This is a common assumption made by prior works [26], and such an API is also provided by OpenAI ($\mathrm{top\_logprobs=5}$), which can benefit the normal users in understanding the model confidence, debugging and analyzing the model’s behavior, customizing sampling strategies, etc. Consider that the top $L=5$ tokens and their probabilities are returned to the attackers. The probability that the attacker can find an unwatermarked token in the token candidates’ list of length $L$ is $1-\gamma^{L}$ for KGW and Unigram, which becomes sufficiently large given $L=5$ and $\gamma=0.5$. The attacker will query the detection using these replacements and sample a token based on their probabilities and detection scores to remove the watermark while preserving a high output quality.
Spoofing Attack. Spoofing attacks follow a similar procedure where the attacker can generate (harmful) content using a local model. When sampling the tokens, instead of selecting those that yield low confidence scores as in removal attacks, the attacker will choose tokens that have higher confidence scores upon watermark detection queries. Thanks to the robustness of the LLM watermarks, attackers don’t need to ensure every single token carries a watermark; only that the overall detection confidence score surpasses the threshold, thereby treating synthesized content as if generated by the watermarked LLM.
(a) Z-Score/P-Value of wm-removal.
(b) Perplexity of wm-removal.
@@ 6.2 Evaluation
Experiment Setup.  We use the same evaluation setup as in Sec. 4.1 and Sec. 5.1. We evaluate the detection scores for both the watermark-removal and the spoofing attacks. We also report the number of queries to the detection API. Furthermore, for the watermark-removal attack, where the attackers care more about the output quality, we report the output PPL. For spoofing attacks, the attackers’ local models are LLAMA-2-7B and OPT-1.3B.
@@
Evaluation Result.  As shown in Fig. 3(a) and Fig. 3(b), watermark-removal attacks exploiting the detection API significantly reduce detection confidence while maintaining high output quality. For instance, for the KGW watermark on LLAMA-2-7B model, we achieve a median z-score of $1.43$, which is much lower than the threshold $4$. The PPL is also close to the watermarked outputs ($6.17$ vs. $6.28$). We observe that the Exp watermark has higher PPL than the other two watermarks. This is because that Exp watermark is deterministic, while other watermarks enable random sampling during inference. Our attack also employs sampling based on the token probabilities and detection scores, thus we can improve the output quality for the Exp watermark.
The spoofing attacks also significantly boost the detection confidence even though the content is not from the watermarked LLM, as depicted in Fig. 3(c). We report the attack success rate (ASR) and the number of queries for both of the attacks in Table 2. The ASR quantifies how much of the generated content surpasses or falls short of the detection threshold. These attacks use a reasonable number of queries to the detection API and achieve high success rate, demonstrating practical feasibility. We observe consistent results on OPT-1.3B, please see Appendix K.